OpenStudy (anonymous):

It takes Jake 9 hours longer than Joe to rebuild an engine. Working together, they can do the job in 20 hours. How long would it take each, working alone, to rebuild an engine?

OpenStudy (anonymous):

take it suzi

OpenStudy (anonymous):

9 = 3^2 20 = 2^2 * 5 3^2 * 2^2 * 5 180 hours?

OpenStudy (anonymous):

jake does it in x hours. Therefore, his rate is 1/x Joe does it in y hours. Therefore, his rate is 1/y 20/y+20/x = 1 y+9 = x 20/y + 20/(y+9) = 1 40y+180= y^2 +9y y^2 -31y +180 = 0 then use the quadratic formula :D